Determine the heat loss of seawater flowing through a 1 inch schedule 40 PVC pipe. The inner diameter of the pipe is 1 inch and the outer diameter is 1.049 inches. The temperature of the water is 50 degrees F and the outside air is 72 degrees F. The seawater is flowing at 3 GPM. The thermal conductivity of the PVC pipe is 0.20 [W/(mK)].  The heat transfer coefficient between the water and the pipe is 75 [W/m2K].

Take the same values and calculate it for schedule 40 10 inch inner diameter and 10.312 inch outer diameter.

Please show each step

After the heat loss is found convert it to degrees per minute.
Answer #1

Date -LI Page No.: | Solutions Too&727 Tw=50F Givent Too=72F = 295.3724 Tw=50F = 283.15% MAAA Now Resistance in heat transfe10.0133 0127 ,025 Raz in (A) In/0.033) 2TTK L 203.148 0.280.0254 (Rz = 1.4456] Rez neroze = 75*°2x3-14 20.0133 48.0254 {Rs =Date -hh Page No.: R₂ = 1 ho 2TRL 7542 403-1480113095700254 R3=0.6388 Now, (295.37 -28.3.15 0:16456442.49 +0:6388 102 = 2.25